OpenStudy (anonymous):

is 4y=5x+2 and 8x=3-10y perpendicular @jim_thompson5910

OpenStudy (anonymous):

Well how do you determine whether two lines are perpendicular or not?

OpenStudy (anonymous):

what does that mean exactly

OpenStudy (anonymous):

when 2 lines are perpendicular is when they form a 90º angle when they touch 

OpenStudy (anonymous):

so that means that they are right

OpenStudy (anonymous):

you would have to see, because that is the condition for them to be perpendicular, you need to put both functions in y=mx+b form so you can determine whether they are or not. it depends solely in the slope of the lines you are analyzing

OpenStudy (anonymous):

The slopes would be opposite reciprocals.

OpenStudy (anonymous):

so first you have to transform your 4y=5x+2 to y=mx+b, that's quite easy, you just need to divide both sides by 4 (you may have learnt this as "passing" the 4 multiplying the y dividing to the other side)

OpenStudy (anonymous):

so it's as simple as \[4y=5x+2,(4y)/4=\frac{ 5x+2 }{ 4 }, y=\frac{ 5 }{ 4 }x+\frac{ 2 }{ 4 }\]

OpenStudy (anonymous):

in the formula y=mx+b, the slope is "m", in this case it would be 5/4

OpenStudy (anonymous):

to determine whether or not 2 lines are perpendicular by using only their formulas, the slope of one should be the opposite by multiplication of the other and with the opposite sign. What this means is that if we have a line of any slope (lets call it "a") the condition that MUST be covered in order for any other line to be perpendicular is that its slope (lets call it b) must be: \[b=-a ^{-1}=-\frac{ 1 }{ a }\]

OpenStudy (anonymous):

so lets get the other slope shall we

OpenStudy (anonymous):

our first slope is 5/4, we have to get the second one from \[8x=3-10y, (8x)-3=(-10y+3)-3, 8x-3=-10y,y=-\frac{ 8x-3 }{ 10 }=-\frac{ 8 }{ 10 }x+\frac{ 3 }{ 10 }\]

OpenStudy (anonymous):

\[y=-\frac{ 8 }{ 10 }x+\frac{ 3 }{ 10 }=-\frac{ 4 }{ 5 }x+\frac{ 3 }{ 10 }\]

OpenStudy (anonymous):

so, now we have the 2 slopes, and we know that fot them to be perpendicular the condition \[b=-\frac{ 1 }{ a }\] should be complete, so our "a"= 5/4, and our b=-4/5\[-\frac{ 1 }{ a }=-\frac{ 1 }{ \frac{ 5 }{ 4 } }\]

OpenStudy (anonymous):

that should be "b" right? can you tell me if it is?

OpenStudy (anonymous):

no ??

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

do you know how to divide fractions?

OpenStudy (anonymous):

yea but I have a lot of trouble with them

OpenStudy (anonymous):

try it out, I know you can do it

OpenStudy (anonymous):

okay I will try but what do I divdie

OpenStudy (anonymous):

you have to divide \[-\frac{ 1/1 }{ 5/4 }\]

OpenStudy (anonymous):

okay I got 0.05 am I ritte

OpenStudy (anonymous):

-0.05

OpenStudy (anonymous):

no, remember, when you are dividing fractions you do this \[\frac{ a }{ b }\div \frac{ c }{ d }=\frac{ ad }{ bc }\]

OpenStudy (anonymous):

how would you do that if we have a=1 b=1 c=5 d=4?

OpenStudy (anonymous):

1/1 divided by 5/4 = 15/14 ritte

OpenStudy (anonymous):

no, ab is short for a*b, its a multiplication

OpenStudy (anonymous):

it's easy, I know you can do it

OpenStudy (anonymous):

1/20 because 1*1=1 and then 5*4=20

OpenStudy (anonymous):

@SalvadorV

OpenStudy (anonymous):

check again \[\frac{ a }{ b }\div \frac{ c }{ d }=\frac{ ad }{ bc }, \frac{ 1 }{ 1 }\div \frac{ 5 }{ 4 }=?\]

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

mmmm nope, check those letters again, see this 

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

buut I still don't noe if they are perpendicular

OpenStudy (anonymous):

ummmm hello @SalvadorV

OpenStudy (anonymous):

remember which were the conditions I told you for them to be perpendicular?

OpenStudy (anonymous):

so yes they are

OpenStudy (anonymous):

sorry, I'm part timing explaining calculus to someone

OpenStudy (anonymous):

remember it was that if b=-(1/a) then it was perpendicular

OpenStudy (anonymous):

so b was -4/5

OpenStudy (anonymous):

and a was 5/4

OpenStudy (anonymous):

so if \[-\frac{ 1 }{ a }\] is the same as "b", then the lines are perpendicular

OpenStudy (anonymous):

can you tell me if they are?

OpenStudy (anonymous):

yess they are !! ritte

OpenStudy (anonymous):

YES!!!!!!!! :D

OpenStudy (anonymous):

yay thanks let me noe when yhu are done with calculus

OpenStudy (anonymous):

I don't think I'll be done soon :/ it's a very wide subject and the person I'm helping doesn't quite understand it